# V Holešovičkách 2, 18000 Praha 8, Czech Republic

### v1

@@ Abstract
Gravitational lensing: strong – Galaxies: clusters: general – Galaxies: halos – Cosmology: dark matter
@@ 1 Introduction
The individual halos, subhalos, or different matter components are generally modeled by simple density or surface-density profiles. In practice, the range of possible profiles is narrowed down by the requirement of an analytic expression for the light-deflection angle. Without it the angle has to be integrated numerically over the surface-density distribution for each light ray, which is computationally prohibitive. As a result, density profiles that cannot be integrated analytically are shunned, even though they may provide a better description of the mass distribution of the lens. Analytic expressions for the deflection angle are available for a number of spherically symmetric lens models (Keeton, 2001). However, in the case of ellipsoidal models, one geometric step closer to reality, analytic formulae for the deflection-angle components have been obtained only for a few.
@@
In this work we derive analytic expressions for the deflection angle of the ellipsoidal NFW model and describe its lensing properties. We start in Sect. 2 from a triaxial homoeoidally symmetric mass distribution with an arbitrary spatial orientation and compute the properties of its corresponding elliptically symmetric surface-density distribution. In Sect. 3 we review the complex formalism of Bourassa & Kantowski (1975) for computing the deflection angle of a lens with an elliptically symmetric surface-density distribution. In Sect. 4 we apply the methodology introduced in the previous sections to the ellipsoidal NFW lens, describing: the convergence profile (Sect. 4.1); the complex scattering function and the deflection angle (Sect. 4.2); the shear and phase (Sect. 4.3) with particular attention to the halo center (Sect. 4.3.2) and the zero-shear points (Sect. 4.3.3); the critical curves and caustics and their variation in the lens-parameter space (Sect. 4.4); the properties of images (Sect. 4.5). We discuss the observational relevance and possible further modifications of the model in Sect. 5, and summarize the main results in Sect. 6. Of the three appendices we would like to draw attention to Appendix C in which we introduce the technique of image-plane analysis of critical curves and caustics, a powerful visual tool for tracking the variations of these curves in the parameter space of the lens.
@@ 3 Deflection angle in the Bourassa & Kantowski formalism
Here $\tilde{x}_{1}$ and $\tilde{x}_{2}$ are coordinates along the major and minor axis of the ellipses, $e$ is their eccentricity, and the semi-major axis $\tilde{a}(\tilde{x}_{1},\tilde{x}_{2})$ is defined in Eq. (6). The upper limit of integration is technically valid for a surface-density distribution extending to infinity. In the case of distributions with a finite extent, it should be replaced by the limiting semi-major axis for positions $\bm{x}$ outside the mass distribution. Note that while it may appear tempting to transfer the first denominator into the square-root expression inside the integral, doing so would cancel natural symmetries of the result such as $I(-\bm{\tilde{}}{x})=-I(\bm{\tilde{}}{x})$ and others (Bray, 1984).
@@ 4.1 Surface density and convergence
Note that for a given triaxial mass distribution in a fixed orientation the entire expression before the square brackets is a constant; only the square brackets and the final parentheses are functions of the plane-of-the-sky position via the semi-major axis $\tilde{a}$. In addition, note that it is advantageous to write the plane-of-the-sky separations and positions in units of the projected scale semi-major axis of the halo,
@@
As a function of orientation, for a given triaxially ellipsoidal halo the square root reaches its maximum value of 1 when viewed down the major axis ($\vartheta=\pi/2$ and $\psi=\pi/2$ or $23\pi/$ 2), its minimum value of $\sqrt{1-e_{2}^{2}}$ down the minor axis ($\vartheta=0$ or $\pi$), and the intermediate value of $\sqrt{1-e_{1}^{2}}$ down the median axis ($\vartheta=\pi/2$ and $\psi=0$ or $\pi$). For a more intuitive understanding of the term, note that the volume of an ellipsoid is proportional to the square root of the numerator, while the area of its elliptical projection is proportional to the square root of the denominator, as indicated by Eq. (74). We may thus re-write Eq. (22) in a form that is easier to interpret,
@@
Figure 2: Lens-plane plots of the deflection angle $\bm{\alpha}_{0}(\bm{x})/\kappa_{\text{s}}$ of the ellipsoidal NFW lens with components given by Eq. (30) and Eq. (31). The three panels correspond to eccentricities $e\in\{0,0.6,0.9\}$, as marked above the plots. The arrows mark the deflection-angle vectors at the position of their central dots. The amplitude of each vector is indicated by its color (with a separate color bar for each panel), and by constant-amplitude contours at values marked vertically across the color bars. Positions with maximum deflection amplitude are marked by the large bold black circle in the left panel, and by the two solid black dots along the minor axis in the second and third panels. The two small black circles in these panels indicate saddle points with maximum deflection along the major axis. Note that in the case of non-zero eccentricity the deflection at positions off the major and minor axes typically has a non-radial orientation.
@@ 4.2 Deflection angle
Figure 3: Left panel: Position of deflection-angle maximum $x_{2\text{max}}$ (dashed line) and position of saddle point $x_{1\text{sad}}$ (solid line) as a function of ellipticity $1-\sqrt{1-e^{2}}$ (linear bottom axis) and eccentricity $e$ (non-linear top axis). Right panel: Deflection-angle amplitudes $|\bm{\alpha}_{0}|$ at these positions in units of convergence parameter $\kappa_{\text{s}}$ as a function of ellipticity and eccentricity.
@@
Both expressions in the curly braces are formed by combinations of the same three transcendental functions defined in Eq. (4.2), each weighted by a polynomial in terms of $x_{i}$. The formulae for the two components thus differ only in these polynomials. Note also that $x_{1}$ cannot be factored from $\alpha_{01}$, just as $x_{2}$ cannot be factored from $\alpha_{02}$ — in both cases due to the $\arctan$ term. This implies that the general orientation of $\bm{\alpha}$ is not radial. An analysis of Eqs. (30) and (31) shows that the deflection is oriented radially only along the major and minor axes.
@@ 4.3.1 Large-scale properties
Lens-plane maps of the shear components, the shear, and the phase are presented in Fig. 4 for projected eccentricities $e=0$, 0.6, and 0.9.
In the zero-eccentricity case shown in the left column, the shear component $\gamma_{1}$ (top row) is zero along both diagonals, and it reaches its azimuthal positive maxima and negative minima along the vertical and horizontal axes, respectively. The component $\gamma_{2}$ (second row) has the same pattern rotated counter-clockwise by $\pi/4$: it is zero along both axes, it reaches its azimuthal extremes along the diagonals, and it alternates in the plot quadrants between negative (first and third) and positive values (second and fourth). The shear (third row) has a purely radial dependence, reaching its maximum value $\gamma=\kappa_{\text{s}}$ at the halo center and decreasing outward asymptotically to zero (e.g., Karamazov & Heyrovský, 2022). The phase $\phi$ (bottom row) has a purely azimuthal dependence: it is equal to the azimuthal angle measured from the horizontal axis shifted by $-\pi/2$ and limited here to the interval $(-\pi/2,\pi/2]$. The phase is thus zero along the vertical axis, $\pi/2$ along the horizontal axis, and it alternates in the plot quadrants between negative and positive values in the same manner as $\gamma_{2}$.
For non-zero eccentricities shown in the two right columns, the character of the plots changes significantly. Least so in the case of $\gamma_{2}$ (second row): the azimuthal extremes occur along lines closer to the major axis, and the absolute values in units of $\kappa_{\text{s}}$ decrease, as indicated by the shrinking contours. In the case of $\gamma_{1}$ (top row), the absolute values decrease in the positive areas along the minor axis, but they increase in the negative areas along the major axis. However, while the inner negative contours of $\gamma_{1}$ expand, the outer ones shrink. Note also that while the zero contours of $\gamma_{2}$ lie always along the coordinate axes irrespective of the eccentricity, the zero-contours of $\gamma_{1}$ do not reach the halo center when $e>0$. The pattern of the shear $\gamma$ (third row) changes dramatically: the outer oval contours give way to the central dipole-like pattern dominated by the character of $\gamma_{1}$. Along the minor axis just above and below the center there are localized low-shear regions. The phase (bottom row) preserves the large-scale pattern in the quadrants, but the contours are tilted toward the horizontal axis. More specifically, the $\gamma_{2}=0$ contours correspond to $\phi=0$ and $\phi=\pi/2$ contours in all the phase plots in Fig. 4 (as well as in Fig. 5), while the $\gamma_{1}=0$ contours correspond to $|\phi|=\pi/4$ contours. Note in the bottom right panel of Fig. 4 that non-zero eccentricities lead to deviations from the large-scale pattern in the vicinity of the halo center, which are explored in the following Sect. 4.3.2.
@@ 4.3.2 Halo-center features
where the expression “$x^{2}\,\ln{x}$” in the next higher order stands for both $x_{i}^{2}\,\ln{a}$ and $x_{i}^{2}\,\ln{x}$ terms. The phase has to be computed from these approximations using Eq. (33) and Eq. (34).
@@
In the bottom row, note that the phase $\phi$ switches along the minor axis from 00 to $\pi/2$ between the zero-shear points. In the immediate vicinity of the center the phase is thus vertical along both axes and near-vertical off the axes. All contours meet at the zero-shear points, with $\phi=0$ pointing outward along the minor axis, $\phi=\pi/2$ inward along the minor axis and all other values along contours pointing sideways under different angles. Contours pointing in opposite directions have perpendicular phases (±$\pm\pi/2$ difference). We conclude that in the case of non-zero eccentricity the phase is defined at the halo center, but it is undefined at the zero-shear points.
@@ 4.3.3 Zero-shear points
As shown in Karamazov & Heyrovský (2022), zero-shear points mark the lens-plane positions of undistorted positive-parity images, while undistorted negative-parity images occur along the unit-convergence contour. In the special case when the zero-shear points lie on the unit-convergence contour, they form umbilic points belonging to the critical curve (see Sect. 4.4). In the case of our lens model the zero-shear points occur for any non-zero eccentricity at the intersections of the $\gamma_{1}=0$ and $\gamma_{2}=0$ contours. The pattern of these contours in Fig.s 4 and 5 shows that the zero-shear points are located at $\bm{x}_{0\,\displaystyle{\gamma}}=(0,\pm\,x_{0\,\displaystyle{\gamma}})$, where $x_{0\,\displaystyle{\gamma}}$ is the positive root of $\gamma_{1}(0,x_{0\,\displaystyle{\gamma}})=0$. Using Eq. (35), this condition can be simplified to
@@ 4.4 Critical curves and caustics
In order to compute the lens-equation Jacobian we combine the expressions for the convergence from Eq. (21) and for the shear components from Eq.s (35) and (36) in the general formula
@@ 4.4.1 Variation of geometries with eccentricity
The tangential critical curve for $e=0$ is the outer circular component. At $e=0.3$ it is oval-shaped with its major axis horizontal, along the major axis of the projected mass distribution. At $e=0.6$ the oval is more compressed, locally concave in the vicinity of the vertical axis. Eccentricity $e=0.811$ corresponds to the hyperbolic-umbilic configuration for $\kappa_{\text{s}}=0.25$. The tangential and radial critical curves come into contact at the two umbilic points along the vertical axis. These are defined by zero shear and unit convergence, as indicated by Eq. (51) and Eq. (52). Note that the tangential critical curve has sharp points with an undefined tangent at the umbilic points, but no inflection points. At $e=0.9$ the tangential critical curve is peanut-shaped, very elongated horizontally and locally concave close to the vertical axis.
@@
The color-map background of the critical-curve plots shows the Jacobian, with shades of red and blue indicating positive and negative values, respectively. Since image magnification is given by the absolute inverse of the Jacobian, paler regions correspond to higher magnification. Note that while these are fairly extended at low eccentricity, at high eccentricity (and the same $\kappa_{\text{s}}$) high magnification is limited to a narrow region along the critical curve.
@@
Figure 8: Low- and intermediate-$\kappa_{\text{s}}$ lensing regimes of the ellipsoidal NFW lens mapped in its parameter space, as described in Sect. 4.4.2 and Sect. 4.4.3. Eccentricity $e$ is marked on the right vertical axes; ellipticity $1-\sqrt{1-e^{2}}$ on the left vertical axes; convergence parameter on the horizontal axis up to $\kappa_{\text{s}}=2.1$ in the left panel and up to $\kappa_{\text{s}}=8$ in the right panel. Plotted boundaries mark the following transitions: cusp piercing (bold solid black); tangential peanut (thin dashed green); radial peanut (thin solid purple); hyperbolic umbilic (dashed black); tangential bulge (thin dotted green); butterfly metamorphosis (bold dashed green). Parameter combinations of the critical-curve and caustic examples presented in Fig. 7 and Fig. 9 are marked by red and blue dots, respectively. Cyan dotted circles mark combinations for which image-orientation plots are presented in Fig. 17; the green dot marks the branching point of the tangential-bulge boundary from the tangential-peanut boundary. Vertical dashed lines: $\kappa_{\text{s}}$ values for which critical-curve and caustic dimensions are plotted in Fig. 12 and Fig. 14, respectively. Horizontal dashed lines: $e$ values for which the image-plane analysis is demonstrated in Fig. 19 and Fig. 20.
4.4.2 Lensing regimes in parameter space: low convergence parameter

### v2

@@ Abstract
gravitational lensing: strong – galaxies: clusters: general – galaxies: halos – dark matter
@@ 1 Introduction
The individual halos, subhalos, or different matter components are generally modeled by simple density or surface-density profiles (for a theoretical overview, see Wagner, 2020). In practice, the range of possible profiles is narrowed down by the requirement of an analytic expression for the light-deflection angle. Without it the angle has to be integrated numerically over the surface-density distribution for each light ray, which is computationally prohibitive. As a result, density profiles that cannot be integrated analytically are shunned, even though they may provide a better description of the mass distribution of the lens. Analytic expressions for the deflection angle are available for a number of spherically symmetric lens models (Keeton, 2001). However, in the case of ellipsoidal models, one geometric step closer to reality, analytic formulae for the deflection-angle components have been obtained only for a few.
@@
In this work we derive analytic expressions for the deflection angle of the ellipsoidal NFW model and describe its lensing properties. We start in Sect. 2 from a triaxial homoeoidally symmetric mass distribution with an arbitrary spatial orientation and compute the properties of its corresponding elliptically symmetric surface-density distribution. In Sect. 3 we review the complex formalism of Bourassa & Kantowski (1975) for computing the deflection angle of a lens with an elliptically symmetric surface-density distribution. In Sect. 4 we apply the methodology introduced in the previous sections to the ellipsoidal NFW lens, describing: the convergence profile (Sect. 4.1); the complex scattering function and the deflection angle (Sect. 4.2); the shear and phase (Sect. 4.3) with particular attention to the halo center (Sect. 4.3.2) and the zero-shear points (Sect. 4.3.3); the critical curves and caustics and their variation in the lens-parameter space (Sect. 4.4); the properties of images (Sect. 4.5). We discuss the observational relevance and possible further modifications of the model in Sect. 5, and summarize the main results in Sect. 6. Of the four appendices we would like to draw attention to Appendix C in which we introduce the technique of image-plane analysis of critical curves and caustics, a powerful visual tool for tracking the variations of these curves in the parameter space of the lens.
@@ 3 Deflection angle in the Bourassa & Kantowski formalism
Here $\tilde{x}_{1}$ and $\tilde{x}_{2}$ are coordinates along the major and minor axis of the ellipses, $e$ is their eccentricity, and the semi-major axis $\tilde{a}(\tilde{x}_{1},\tilde{x}_{2})$ is defined in Eq. (6). The upper limit of integration is technically valid for a surface-density distribution extending to infinity. In the case of distributions with a finite extent, it should be replaced by the limiting semi-major axis for positions $\bm{x}$ outside the mass distribution. We note that while it may appear tempting to transfer the first denominator into the square-root expression inside the integral, doing so would cancel natural symmetries of the result such as $I(-\bm{\tilde{}}{x})=-I(\bm{\tilde{}}{x})$ and others (Bray, 1984).
@@ 4.1 Surface density and convergence
We note that for a given triaxial mass distribution in a fixed orientation the entire expression before the square brackets is a constant; only the square brackets and the final parentheses are functions of the plane-of-the-sky position via the semi-major axis $\tilde{a}$. In addition, we note that it is advantageous to write the plane-of-the-sky separations and positions in units of the projected scale semi-major axis of the halo,
@@
As a function of orientation, for a given triaxially ellipsoidal halo the square root reaches its maximum value of 1 when viewed down the major axis ($\vartheta=\pi/2$ and $\psi=\pi/2$ or $23\pi/$ 2), its minimum value of $\sqrt{1-e_{2}^{2}}$ down the minor axis ($\vartheta=0$ or $\pi$), and the intermediate value of $\sqrt{1-e_{1}^{2}}$ down the median axis ($\vartheta=\pi/2$ and $\psi=0$ or $\pi$). For a more intuitive understanding of the term, we note that the volume of an ellipsoid is proportional to the square root of the numerator, while the area of its elliptical projection is proportional to the square root of the denominator, as indicated by Eq. (74). We may thus re-write Eq. (22) in a form that is easier to interpret,
@@
Figure 2: Lens-plane plots of the deflection angle $\bm{\alpha}_{0}(\bm{x})/\kappa_{\text{s}}$ of the ellipsoidal NFW lens with components given by Eqs. (30) and (31). The three panels correspond to eccentricities $e\in\{0,0.6,0.9\}$, as marked above the plots. The arrows mark the deflection-angle vectors at the position of their central dots. The amplitude of each vector is indicated by its color (with a separate color bar for each panel), and by constant-amplitude contours at values marked vertically across the color bars. Positions with maximum deflection amplitude are marked by the large bold black circle in the left panel, and by the two solid black dots along the minor axis in the second and third panels. The two small black circles in these panels indicate saddle points with maximum deflection along the major axis. We note that in the case of non-zero eccentricity the deflection at positions off the major and minor axes has a non-radial orientation.
@@ 4.2 Deflection angle
Figure 3: Scaling of the deflection-angle maps with eccentricity. Left panel: Position of deflection-angle maximum $x_{2\text{max}}$ (dashed line) and position of saddle point $x_{1\text{sad}}$ (solid line) as a function of eccentricity $e$ (non-linear top axis) and ellipticity $1-\sqrt{1-e^{2}}$ (linear bottom axis). Right panel: Deflection-angle amplitudes $|\bm{\alpha}_{0}|$ at these positions in units of convergence parameter $\kappa_{\text{s}}$ as a function of eccentricity and ellipticity.
@@
Both expressions in the curly braces are formed by combinations of the same three transcendental functions defined in Eq. (4.2), each weighted by a polynomial in terms of $x_{i}$. The formulae for the two components thus differ only in these polynomials. We note also that $x_{1}$ cannot be factored from $\alpha_{01}$, just as $x_{2}$ cannot be factored from $\alpha_{02}$ — in both cases due to the $\arctan$ term. This implies that the general orientation of $\bm{\alpha}$ is not radial. An analysis of Eqs. (30) and (31) shows that the deflection is oriented radially only along the major and minor axes.
@@ 4.3.1 Large-scale properties
Lens-plane maps of the shear components, the shear, and the phase are presented in Fig. 4 for projected eccentricities $e=0$, 0.6, and 0.9. In the zero-eccentricity case shown in the left column, the shear component $\gamma_{1}$ (top row) is zero along both diagonals, and it reaches its azimuthal positive maxima and negative minima along the vertical and horizontal axes, respectively. The component $\gamma_{2}$ (second row) has the same pattern rotated counter-clockwise by $\pi/4$: it is zero along both axes, it reaches its azimuthal extremes along the diagonals, and it alternates in the plot quadrants between negative (first and third) and positive values (second and fourth). The shear (third row) has a purely radial dependence, reaching its maximum value $\gamma=\kappa_{\text{s}}$ at the halo center and decreasing outward asymptotically to zero (e.g., Karamazov & Heyrovský, 2022). The phase $\phi$ (bottom row) has a purely azimuthal dependence: it is equal to the azimuthal angle measured from the horizontal axis shifted by $-\pi/2$ and limited here to the interval $(-\pi/2,\pi/2]$. The phase is thus zero along the vertical axis, $\pi/2$ along the horizontal axis, and it alternates in the plot quadrants between negative and positive values in the same manner as $\gamma_{2}$.
For non-zero eccentricities shown in the two right columns, the character of the plots changes significantly. Least so in the case of $\gamma_{2}$ (second row): the azimuthal extremes occur along lines closer to the major axis, and the absolute values in units of $\kappa_{\text{s}}$ decrease, as indicated by the shrinking contours. In the case of $\gamma_{1}$ (top row), the absolute values decrease in the positive areas along the minor axis, but they increase in the negative areas along the major axis. However, while the inner negative contours of $\gamma_{1}$ expand, the outer ones shrink. We note also that while the zero contours of $\gamma_{2}$ lie always along the coordinate axes irrespective of the eccentricity, the zero-contours of $\gamma_{1}$ do not reach the halo center when $e>0$. The pattern of the shear $\gamma$ (third row) changes dramatically: the outer oval contours give way to the central dipole-like pattern dominated by the character of $\gamma_{1}$. Along the minor axis just above and below the center there are localized low-shear regions. The phase (bottom row) preserves the large-scale pattern in the quadrants, but the contours are tilted toward the horizontal axis. More specifically, the $\gamma_{2}=0$ contours correspond to $\phi=0$ and $\phi=\pi/2$ contours in all the phase plots in Fig. 4 (as well as in Fig. 5), while the $\gamma_{1}=0$ contours correspond to $|\phi|=\pi/4$ contours. It can be seen in the bottom right panel of Fig. 4 that non-zero eccentricities lead to deviations from the large-scale pattern in the vicinity of the halo center, which are explored in the following Sect. 4.3.2.
@@ 4.3.2 Halo-center features
where the expression “$x^{2}\,\ln{x}$” in the next higher order stands for both $x_{i}^{2}\,\ln{a}$ and $x_{i}^{2}\,\ln{x}$ terms. The phase has to be computed from these approximations using Eqs. (33) and (34).
@@
In the bottom row, we note that the phase $\phi$ switches along the minor axis from 00 to $\pi/2$ between the zero-shear points. In the immediate vicinity of the center the phase is thus vertical along both axes and near-vertical off the axes. All contours meet at the zero-shear points, with $\phi=0$ pointing outward along the minor axis, $\phi=\pi/2$ inward along the minor axis and all other values along contours pointing sideways under different angles. Contours pointing in opposite directions have perpendicular phases (±$\pm\pi/2$ difference). We conclude that in the case of non-zero eccentricity the phase is defined at the halo center, but it is undefined at the zero-shear points.
@@ 4.3.3 Zero-shear points
As shown in Karamazov & Heyrovský (2022), zero-shear points mark the lens-plane positions of undistorted positive-parity images, while undistorted negative-parity images occur along the unit-convergence contour. In the special case when the zero-shear points lie on the unit-convergence contour, they form umbilic points belonging to the critical curve (see Sect. 4.4). In the case of our lens model the zero-shear points occur for any non-zero eccentricity at the intersections of the $\gamma_{1}=0$ and $\gamma_{2}=0$ contours. The pattern of these contours in Figs. 4 and 5 shows that the zero-shear points are located at $\bm{x}_{0\,\displaystyle{\gamma}}=(0,\pm\,x_{0\,\displaystyle{\gamma}})$, where $x_{0\,\displaystyle{\gamma}}$ is the positive root of $\gamma_{1}(0,x_{0\,\displaystyle{\gamma}})=0$. Using Eq. (35), this condition can be simplified to
@@ 4.4 Critical curves and caustics
In order to compute the lens-equation Jacobian we combine the expressions for the convergence from Eq. (21) and for the shear components from Eqs. (35) and (36) in the general formula
@@ 4.4.1 Variation of geometries with eccentricity
The tangential critical curve for $e=0$ is the outer circular component. At $e=0.3$ it is oval-shaped with its major axis horizontal, along the major axis of the projected mass distribution. At $e=0.6$ the oval is more compressed, locally concave in the vicinity of the vertical axis. Eccentricity $e=0.811$ corresponds to the hyperbolic-umbilic configuration for $\kappa_{\text{s}}=0.25$. The tangential and radial critical curves come into contact at the two umbilic points along the vertical axis. These are defined by zero shear and unit convergence, as indicated by Eqs. (51) and (52). We note that the tangential critical curve has sharp points with an undefined tangent at the umbilic points, but no inflection points. At $e=0.9$ the tangential critical curve is peanut-shaped, very elongated horizontally and locally concave close to the vertical axis.
@@
The color-map background of the critical-curve plots shows the Jacobian, with shades of red and blue indicating positive and negative values, respectively. Since image magnification is given by the absolute inverse of the Jacobian, paler regions correspond to higher magnification. We note that while these are fairly extended at low eccentricity, at high eccentricity (and the same $\kappa_{\text{s}}$) high magnification is limited to a narrow region along the critical curve.
@@
Figure 8: Low- and intermediate-$\kappa_{\text{s}}$ lensing regimes of the ellipsoidal NFW lens mapped in its parameter space, as described in Sect. 4.4.2 and Sect. 4.4.3. Eccentricity $e$ is marked on the right vertical axes; ellipticity $1-\sqrt{1-e^{2}}$ on the left vertical axes; convergence parameter on the horizontal axis up to $\kappa_{\text{s}}=2.1$ in the left panel and up to $\kappa_{\text{s}}=8$ in the right panel. Plotted boundaries mark the following transitions: cusp piercing (bold solid black); tangential peanut (thin dashed green); radial peanut (thin solid purple); hyperbolic umbilic (dashed black); tangential bulge (thin dotted green); butterfly metamorphosis (bold dashed green). Parameter combinations of the critical-curve and caustic examples presented in Figs. 7 and 9 are marked by red and blue dots, respectively. Cyan dotted circles mark combinations for which image-orientation plots are presented in Fig. 16; the green dot marks the branching point of the tangential-bulge boundary from the tangential-peanut boundary. Vertical dashed lines: $\kappa_{\text{s}}$ values for which critical-curve and caustic dimensions are plotted in Figs. 11 and 13, respectively. Horizontal dashed lines: $e$ values for which the image-plane analysis is demonstrated in Figs. 18 and 19.
4.4.2 Lensing regimes in parameter space at low $\kappa_{\text{s}}$